# ConDiff: A Challenging Dataset for Neural Solvers of Partial Differential Equations

### v1

@@ Abstract
We present ConDiff, a novel dataset for scientific machine learning. ConDiff focuses on the diffusion equation with varying coefficients, a fundamental problem in many applications of parametric partial differential equations (PDEs). The main novelty of the proposed dataset is that we consider discontinuous coefficients with high contrast. These coefficient functions are sampled from a selected set of distributions. This class of problems is not only of great academic interest, but is also the basis for describing various environmental and industrial problems. In this way, ConDiff shortens the gap with real-world problems while remaining fully synthetic and easy to use. ConDiff consists of a diverse set of diffusion equations with coefficients covering a wide range of contrast levels and heterogeneity with a measurable complexity metric for clearer comparison between different coefficient functions. We baseline ConDiff on standard deep learning models in the field of scientific machine learning. By providing a large number of problem instances, each with its own coefficient function and right-hand side, we hope to encourage the development of novel physics-based deep learning approaches, such as neural operators and physics-informed neural networks, ultimately driving progress towards more accurate and efficient solutions of complex PDE problems.
@@ 1 Introduction
In recent years, machine learning techniques have emerged as a promising approach to solving PDEs, offering a new perspective in scientific computing. Machine learning algorithms, especially those based on neural networks, have demonstrated success in approximating complex functions and physical phenomena. Neural networks can provide more efficient and scalable methods compared to traditional numerical methods, which can be computationally expensive and limited by the dimensionality of the problem to be solved. Approaches using physical losses Karniadakis et al. (2021), operator learning Li et al. (2020), symmetries incorporation Wang et al. (2020), data-driven discretization Bar-Sinai et al. (2019) lead to more physically meaningful solutions and gave neural networks better recognition than just black-boxes.
Classical methods for solving PDEs have been extensively developed and refined over the years, providing a basis for understanding and analyzing various physical phenomena. These methods involve discretization the PDEs using techniques as the finite difference method LeVeque (2007), finite element method Bathe (2006), finite volume method Eymard et al. (2000) or spectral methods Trefethen (2000), followed by numerical solution of the resulting algebraic equations. While these methods have been successful in solving a wide range of PDEs, they often face the cure of dimensionality when parametric PDEs need to be solved in connection with optimization, optimal control, parameter identification, uncertainty quantification. The reduction of complexity for such classes of problems can be addressed with surrogate models using machine learning.
The main approaches in scientific machine learning are (i) using governing equations as loss functions with physics-informed neural networks Karniadakis et al. (2021); Cai et al. (2021); Eivazi et al. (2024); Raissi et al. (2019); (ii) learning mappings between infinite-dimensional function spaces with neural operators Li et al. (2020); Fanaskov and Oseledets (2023); Lu et al. (2021a); Li et al. (2024); Tran et al. (2021); (iii) hybrid approaches where machine learning techniques are incorporated into classical simulations Brunton and Kutz (2022); Schnell and Thuerey (2024); Hsieh et al. (2019); Ingraham et al. (2018).
@@
With ConDiff (short for Contrast Diffusion), we propose a novel dataset that complements existing datasets and provides an approach to generating complex coefficients for parametric PDEs that can address real-world problems. We also provide a measurable metric for the complexity of the dataset and validate our approach on a list of well-known models.
@@ 2 ConDiff
Creating a comprehensive benchmark for classes of parametric PDEs is a particular challenge for the scientific machine learning community. The main challenges in creating a comprehensive dataset are: (i) computational complexity; (ii) storage complexity for the desired dimensions of the discretized PDE and parameter space; (iii) properties of the coefficients and solution functions; (iv) relation to real-world problems. The first and second reasons illustrate a technical bottleneck in the creation of the dataset and are mostly dependent on the hardware and efficiency of the numerical method used. Properties such as coefficient smoothness, discontinuity, spatial variation of the coefficients, variance of the parametric space significantly affect the complexity of the dataset and should be carefully chosen. The solution to PDEs (i.e. the ground truth for the dataset) depends on a number of numerical aspects such as choice of mesh, discretization, numerical algorithm, boundary and initial conditions. Therefore, it is very important to consider every little detail regarding different numerical schemes, PDEs, boundary and initial conditions.
Existing benchmarks and datasets cover different aspects of scientific machine learning for different classes of PDEs and can be divided into several groups. PDEBench Takamoto et al. (2022), PINNacle Hao et al. (2023), CFDBench Luo et al. (2023) have a large number of PDEs with different boundary and initial conditions and different dimensionality and resolution. The best covered area is weather forecasting: SuperBench Ren et al. (2023), ClimSim Yu et al. (2024), DynaBench Dulny et al. (2023), OceanBench Johnson et al. (2024), ChaosBench Nathaniel et al. (2024). There are also domain specific datasets with applications to Lagrangian mechanics LagrangeBench Toshev et al. (2024) and phase change phenomena BubbleML Hassan et al. (2023). Worth noting frameworks for differential simulations and general environments for PDEs in scientific machine learning: PDE Control Gym Bhan et al. (2024), PDEArena Gupta and Brandstetter (2022), DiffTaichi Hu et al. (2019), DeepXDE Lu et al. (2021b) and $\Phi_{\text{Flow}}$ Holl et al. (2020).
While all of these datasets contribute significantly to the community, to the best of the authors’ knowledge there is no dataset dedicated to the very important class of academic and real-world problems, the class of parametric PDEs with random coefficients. Typically, when a new model is proposed, authors test it with a set of equations with smooth coefficients Brandstetter et al. (2022); Nguyen et al. (2023); Ripken et al. (2023); Bryutkin et al. (2024). Such coefficients do not allow important classes of industrial applications to be addressed. In section 3 we show that increasing the heterogeneity and contrast in the coefficient function leads to increasing challenges in building accurate surrogate models.
@@
Existing benchmarks Takamoto et al. (2022); Hao et al. (2023); Luo et al. (2023) cover a set of PDEs, both steady-state and time-dependent, with different resolutions and time lengths. In our work, we approach the problem from the other side tacking a fixed parametric PDE and generating a comprehensive set of random coefficients for it. We consider a 2D steady-state diffusion equation:
@@
Note that the equation (1) models not only diffusion, but also steady-state Darcy flow in porous media, steady-state heat conduction, etc. To address certain real-world problems, we use the Gaussian Random Field (GRF) to generate the field $\phi(x)$ with the following covariance models as functions of distance $d$:
@@
In the Table 1 one can observe that increasing $\sigma^{2}$ leads to a higher condition number $\kappa(A)=|\lambda_{\max}|\big{/}|\lambda_{\min}|$ of the discretized differential operator Capizzano (2003). The condition number is closely related to the performance of the numerical methods used to solve PDEs Benzi et al. (2005); Elman et al. (2014). A high condition number indicates that small changes in the input can lead to large changes in the output, making the problem ill-conditioned. This is particularly important in PDEs, where small perturbations can significantly affect the solution. Also, if iterative methods are used to solve the discretized PDE, a larger condition number means a larger number of iterations for unpreconditioned and most of preconditioned iterative methods Saad (2003).
@@
All of the above reasoning is done with regard to the frequent occurrence of such tasks in real world Hashmi (2014); Massimo (2013); Carr and Turner (2016); Oristaglio and Hohmann (1984); Muravleva et al. (2021), including composite materials modeling, heat transfer, geophysical problems, fluid flow modeling. In Figure 2 one can see a cross section of the $x-$permeability field along the $z$ axis over the SPE10 model 2 benchmark Christie and Blunt (2001). The term permeability is used to denote the coefficients of the above equation when considering flow in porous media. This field is very similar to the ConDiff samples in Figure 1.
@@
For each parameter set, we generate $1000$ training and $200$ test realizations of the diffusion equation (1) on $64\times 64$ and $128\times 128$ grids. Note that datasets with the same field parameters but different grid sizes are generated independently and do not represent the same field. The fixed geometry of ConDiff allows PDEs with different fields $\phi(x)$ to be compared without fear that different geometries will interfere with a fair comparison across different coefficient functions. To control the complexity of the generated PDEs realizations, we set contrast bounds during generation as follows:
@@ 3 Experiments
We do not attempt to benchmark every scientific machine learning surrogate model on the ConDiff. We have selected the following list of models to validate on the ConDiff: Spectral Neural Operator (SNO) Fanaskov and Oseledets (2023), Factorized Fourier Neural Operator (F-FNO) Tran et al. (2021), Dilated ResNet (DilResNet) Yu et al. (2017) and U-Net Ronneberger et al. (2015). Neural operators FNO and SNO are both types of neural networks designed to learn mappings between function spaces, in particular to solve PDEs. Neural operators are designed to be universal approximators of continuous operators acting between Banach spaces and to be discretization invariant, meaning that they can handle different discretizations of the underlying function spaces without requiring changes to the model. DilResNet and U-Net are classical neural network models originating from the field of computer vision (CV). Both models have shown their applicability beyond CV and have been used extensively for modeling physical phenomena Stachenfeld et al. (2021); Ma et al. (2021). More details about the models used can be found in the Appendix A.
@@
For training neural networks we use frameworks from the JAX Bradbury et al. (2018) ecosystem: Equinox Kidger and Garcia (2021) and Optax DeepMind et al. (2020). The loss function used is the relative $L_{2}$ loss:
@@
4 Conclusion and Limitations
In summary, we propose a novel open-source dataset for the field of neural solving of PDEs. The unique feature of the dataset is discontinuous coefficients with high contrast for parametric PDEs from different distributions. By designing the coefficients in this way, we achieve a high complexity of the generated PDEs, which also illustrates real-world problems. The proposed complexity function allows to distinguish between the generated PDEs. We also provide code to generate new data based on the approach used in this paper. Furthermore, we validate a number of surrogate models on the ConDiff to illustrate its usefulness in the field of scientific machine learning.
@@ 4 Conclusion and Limitations
    Limited number of models validated on ConDiff. Comprehensive benchmark with different approaches (e.g. physics-informed neural networks, hybrid solvers) is needed.
  • 5.
@@ Appendix A Architectures
    F-FNO – Factorized Fourier Neural Operator (F-FNO) from Tran et al. [2021].
@@
    fSNO – Spectral Neural Operator (SNO). The construction mirrors FNO, but instead of FFT, a transformation based on Gauss quadratures is used Fanaskov and Oseledets [2023].
@@
    DilResNet – Dilated Residual Network from Yu et al. [2017], Stachenfeld et al. [2021].
@@
    U-Net – classical computer vision architecture introduced in Ronneberger et al. [2015].
A.1 F-FNO
Unlike the original Li et al. [2020], the authors of Tran et al. [2021] proposed to changing the operator layer to:
@@ A.1 F-FNO
A.2 SNO
We utilized spectral neural operators (SNO) Fanaskov and Oseledets [2023] with linear integral kernels:
@@ A.2 SNO
These linear integral kernels are an extension of the integral kernels used in the FNO Li et al. [2020]. More specifically, starting from the input function $u^{n}$, we produce the output function $u^{n+1}$, which is later transformed by nonlinear activation. The transformation depends on the set of polynomials $p_{j}$ that form a suitable basis for the problem at hand (e.g. trigonometric polynomials, Chebyshev polynomials, etc.). These polynomials are chosen beforehand and do not change during training. The transformation is naturally divided into three parts: analysis, processing, synthesis.
@@
$u^{n+1}=\sum_{j}p_{j}\alpha_{j}^{\textquoteleft}.$
We use SNO in Fourier basis (see Fanaskov and Oseledets [2023]) with encoder-processor-decoder architecture. The number of SNO layers is $4$ and the number of $p_{j}(x)$ is $20$. We use GeLU as activation function.
A.3 DilResNet
The conventional dilated residual network was first proposed in Stachenfeld et al. [2021]. In this study, the DilResNet architecture is configured with four blocks, each consisting of a sequence of convolutions with steps of $[1,2,4,8,4,2,1]$ and a kernel size of $3$. Skip connections are also applied after each block and the GeLU activation function is used.
A.4 U-Net
We adopt the traditional U-Net architecture proposed in Ronneberger et al. [2015]. This U-Net configuration is characterised by a series of levels, where each level has approximately half the resolution of the previous one, and the number of features is doubled. At each level, we apply a sequence of three convolutions, followed by max pooling, and then a transposed convolution for upsampling. After upsampling, three more convolutions are applied at each level. The U-Net used in this study consists of four layers and incorporates the GeLU activation function.

### v2

@@ Abstract
We present ConDiff, a novel dataset for scientific machine learning. ConDiff focuses on the parametric diffusion equation with space dependent coefficients, a fundamental problem in many applications of partial differential equations (PDEs). The main novelty of the proposed dataset is that we consider discontinuous coefficients with high contrast. These coefficient functions are sampled from a selected set of distributions. This class of problems is not only of great academic interest, but is also the basis for describing various environmental and industrial problems. In this way, ConDiff shortens the gap with real-world problems while remaining fully synthetic and easy to use. ConDiff consists of a diverse set of diffusion equations with coefficients covering a wide range of contrast levels and heterogeneity with a measurable complexity metric for clearer comparison between different coefficient functions. We baseline ConDiff on standard deep learning models in the field of scientific machine learning. By providing a large number of problem instances, each with its own coefficient function and right-hand side, we hope to encourage the development of novel physics-based deep learning approaches, such as neural operators, ultimately driving progress towards more accurate and efficient solutions of complex PDE problems.
@@ 1 Introduction
In recent years, machine learning techniques have emerged as a promising approach to solving PDEs, offering a new perspective in scientific computing. Machine learning algorithms, especially those based on neural networks, have demonstrated success in approximating complex functions and physical phenomena. Neural networks can provide more efficient and scalable methods compared to traditional numerical methods, which can be computationally expensive and limited by the dimensionality of the problem to be solved. Approaches using physical losses (Karniadakis et al., 2021), operator learning (Li et al., 2020), symmetries incorporation (Wang et al., 2020), data-driven discretization (Bar-Sinai et al., 2019) lead to more physically meaningful solutions and gave neural networks better recognition than just black-boxes.
Classical methods for solving PDEs have been extensively developed and refined over the years, providing a basis for understanding and analyzing various physical phenomena. These methods involve discretization the PDEs using techniques as the finite difference method (LeVeque, 2007), finite element method (Bathe, 2006), finite volume method (Eymard et al., 2000) or spectral methods (Trefethen, 2000), followed by numerical solution of the resulting algebraic equations. While these methods have been successful in solving a wide range of PDEs, they often face the curse of dimensionality when parametric PDEs need to be solved in connection with optimization, optimal control, parameter identification, uncertainty quantification. The reduction of complexity for such classes of problems can be addressed with surrogate models using machine learning.
The main approaches in scientific machine learning are (i) using governing equations as loss functions with physics-informed neural networks (Karniadakis et al., 2021; Cai et al., 2021; Eivazi et al., 2024; Raissi et al., 2019); (ii) learning mappings between infinite-dimensional function spaces with neural operators (Li et al., 2020; Fanaskov & Oseledets, 2023; Lu et al., 2021a; Li et al., 2024; Tran et al., 2021); (iii) hybrid approaches where machine learning techniques are incorporated into classical simulations (Brunton & Kutz, 2022; Schnell & Thuerey, 2024; Hsieh et al., 2019; Ingraham et al., 2018).
@@
Typically, scientific machine learning datasets have a large number of parametric PDEs (Takamoto et al., 2022; Luo et al., 2023; Hao et al., 2023) that have a single example per PDE. With ConDiff (short for Contrast Diffusion) we focused on the idea of providing a large number of different realizations for a single problem - the diffusion equation. Currently, ConDiff consists of a diverse set of diffusion equations with $24$ realizations, which can be distinguished by complexity, and results in a total of $28800$ samples. We also propose an approach to generating complex coefficients for parametric PDEs that can address real-world problems with a measurable metric of the complexity of the dataset.
@@ 2 ConDiff
Creating a comprehensive benchmark for classes of parametric PDEs is a particular challenge for the scientific machine learning community. The main challenges in creating a comprehensive dataset are: (i) computational complexity; (ii) storage complexity for the desired dimensions of the discretized PDE and parameter space; (iii) properties of the coefficients and solution functions; (iv) relation to real-world problems. The first and second reasons illustrate a technical bottleneck in the creation of the dataset and are mostly dependent on the hardware and efficiency of the numerical method used. Properties such as coefficient smoothness, discontinuity, spatial variation of the coefficients, variance of the parametric space significantly affect the complexity of the dataset and should be carefully chosen. The solution to parametric PDEs (i.e. the ground truth for the dataset) depends on a number of numerical aspects such as choice of mesh, discretization, numerical algorithm, boundary and initial conditions. Therefore, it is very important to consider every little detail regarding different numerical schemes, PDEs, boundary and initial conditions.
Existing benchmarks and datasets cover different aspects of scientific machine learning for different classes of PDEs and can be divided into several groups. PDEBench (Takamoto et al., 2022), PINNacle (Hao et al., 2023), CFDBench (Luo et al., 2023) have a large number of PDEs with different boundary and initial conditions and different dimensionality and resolution. The best covered area is weather forecasting: SuperBench (Ren et al., 2023), ClimSim (Yu et al., 2024), DynaBench (Dulny et al., 2023), OceanBench (Johnson et al., 2024), ChaosBench (Nathaniel et al., 2024). There are also domain specific datasets with applications to Lagrangian mechanics LagrangeBench (Toshev et al., 2024) and phase change phenomena BubbleML (Hassan et al., 2023). Recently, the FlowBench (Tali et al., 2024) dataset with complex geometries was introduced. Worth noting frameworks for differential simulations and general environments for PDEs in scientific machine learning: PDE Control Gym (Bhan et al., 2024), PDEArena (Gupta & Brandstetter, 2022), DiffTaichi (Hu et al., 2019), DeepXDE (Lu et al., 2021b) and $\Phi_{\text{Flow}}$ (Holl et al., 2020).
While all of these datasets contribute significantly to the community, to the best of the authors’ knowledge there is no dataset dedicated to the very important class of academic and real-world problems, the class of parametric PDEs with random coefficients. Typically, when a new model is proposed, authors test it with a set of equations with smooth coefficients (Brandstetter et al., 2022; Nguyen et al., 2023; Ripken et al., 2023; Bryutkin et al., 2024). Such coefficients do not allow important classes of industrial applications to be addressed. In section 3 we show that increasing the heterogeneity and contrast in the coefficient function leads to increasing challenges in building accurate surrogate models.
@@
Existing benchmarks (Takamoto et al., 2022; Hao et al., 2023; Luo et al., 2023) cover a set of PDEs, both steady-state and time-dependent, with different resolutions and time lengths. In our work, we approach the problem from the other side tacking a fixed parametric PDE and generating a comprehensive set of random coefficients for it. We consider a 2D steady-state diffusion equation:
@@
Note that the equation (1) models not only diffusion, but also steady-state Darcy flow in porous media, steady-state heat conduction, etc. To address certain real-world problems, we use the Gaussian Random Field (GRF) to generate the field $\phi(x)$ (Figure 1) with the following covariance models as functions of distance $d$:
@@
In the Table 1 one can observe that increasing $\sigma^{2}$ leads to a higher condition number $\kappa(A)=|\lambda_{\max}|\big{/}|\lambda_{\min}|$ of the discretized differential operator (Capizzano, 2003). The condition number is closely related to the performance of the numerical methods used to solve PDEs (Benzi et al., 2005; Elman et al., 2014). A high condition number indicates that small changes in the input can lead to large changes in the output, making the problem ill-conditioned. This is particularly important in PDEs, where small perturbations can significantly affect the solution. Also, if iterative methods are used to solve the discretized PDE, a larger condition number means a larger number of iterations for unpreconditioned and most of preconditioned iterative methods (Saad, 2003).
@@
All of the above reasoning is done with regard to the frequent occurrence of such tasks in real world (Hashmi, 2014; Massimo, 2013; Carr & Turner, 2016; Oristaglio & Hohmann, 1984; Muravleva et al., 2021), including composite materials modeling, heat transfer, geophysical problems, fluid flow modeling. In Figure 2 one can see a cross section of the $x-$permeability field along the $z$ axis over the SPE10 model 2 benchmark (Christie & Blunt, 2001). The term permeability is used to denote the coefficients of the above equation when considering flow in porous media. This field is very similar to the ConDiff samples in Figure 1.
@@
For each parameter set, we generate $1000$ training and $200$ test realizations of the diffusion equation (1) on $64\times 64$ and $128\times 128$ grids. We provide the train-test split in the ConDiff for fair comparison in future research papers. Note that datasets with the same field parameters but different grid sizes are generated independently and do not represent the same field. The fixed geometry of ConDiff allows PDEs with different fields $\phi(x)$ to be compared without fear that different geometries will interfere with a fair comparison across different coefficient functions. To control the complexity of the generated PDEs realizations, we set contrast bounds during generation as follows:
@@ 3 Experiments
We do not attempt to benchmark every scientific machine learning surrogate model on the ConDiff. Since the ConDiff consists of triplets $\big{(}k(x),f(x),u(x)\big{)}$, its primary use is to validate different architectures of neural operators. Therefore, we have selected the following list of models to validate on the ConDiff: Spectral Neural Operator (SNO) (Fanaskov & Oseledets, 2023), Factorized Fourier Neural Operator (F-FNO) (Tran et al., 2021), Dilated ResNet (DilResNet) (Yu et al., 2017) and U-Net (Ronneberger et al., 2015). Neural operators FNO and SNO are both types of neural networks designed to learn mappings between function spaces, in particular to solve PDEs. Neural operators are designed to be universal approximators of continuous operators acting between Banach spaces and to be discretization invariant, meaning that they can handle different discretizations of the underlying function spaces without requiring changes to the model. DilResNet and U-Net are classical neural network models originating from the field of computer vision (CV). Both models have shown their applicability beyond CV and have been used extensively for modeling physical phenomena (Stachenfeld et al., 2021; Ma et al., 2021). More details about the models used can be found in the Appendix A.1.
@@
For training neural networks we use frameworks from the JAX (Bradbury et al., 2018) ecosystem: Equinox (Kidger & Garcia, 2021) and Optax (DeepMind et al., 2020). The loss function used is the relative $L_{2}$ loss:
@@
4 Discussion
We propose a novel dataset for the field of neural solving of parametric PDEs. The unique feature of the dataset is discontinuous coefficients with high contrast for parametric PDEs from different distributions. By designing the coefficients in this way, we achieve a high complexity of the generated PDEs, which also illustrates real-world problems. The proposed complexity function allows to distinguish between the generated PDEs. We also provide code to generate new data based on the approach used in this paper. Furthermore, we validate a number of surrogate models on the ConDiff to illustrate its usefulness in the field of scientific machine learning.
The practical use of ConDiff is straightforward: it should be used for novel deep learning models and approaches for modeling solution of parametric PDEs from their coefficients. Ultimately, novel deep learning models should exhibit machine-precision prediction quality and not degrade with increasing contrast.
@@ A.1 Architectures
    F-FNO – Factorized Fourier Neural Operator (F-FNO) from (Tran et al., 2021).
@@
    fSNO – Spectral Neural Operator (SNO). The construction mirrors FNO, but instead of FFT, a transformation based on Gauss quadratures is used (Fanaskov & Oseledets, 2023).
@@
    DilResNet – Dilated Residual Network from (Yu et al., 2017), (Stachenfeld et al., 2021).
@@
    U-Net – classical computer vision architecture introduced in (Ronneberger et al., 2015).
F-FNO
Unlike the original (Li et al., 2020), the authors of (Tran et al., 2021) proposed to changing the operator layer to:
@@
SNO
We utilized spectral neural operators (SNO) (Fanaskov & Oseledets, 2023) with linear integral kernels:
@@
These linear integral kernels are an extension of the integral kernels used in the FNO (Li et al., 2020). More specifically, starting from the input function $u^{n}$, we produce the output function $u^{n+1}$, which is later transformed by nonlinear activation. The transformation depends on the set of polynomials $p_{j}$ that form a suitable basis for the problem at hand (e.g. trigonometric polynomials, Chebyshev polynomials, etc.). These polynomials are chosen beforehand and do not change during training. The transformation is naturally divided into three parts: analysis, processing, synthesis.
@@
$u^{n+1}=\sum_{j}p_{j}\alpha_{j}^{‘}.$
We use SNO in Fourier basis (see (Fanaskov & Oseledets, 2023)) with encoder-processor-decoder architecture. The number of SNO layers is $4$ and the number of $p_{j}(x)$ is $20$. We use GeLU as activation function.
DilResNet
The conventional dilated residual network was first proposed in (Stachenfeld et al., 2021). In this study, the DilResNet architecture is configured with four blocks, each consisting of a sequence of convolutions with steps of $[1,2,4,8,4,2,1]$ and a kernel size of $3$. Skip connections are also applied after each block and the GeLU activation function is used.
U-Net
We adopt the traditional U-Net architecture proposed in (Ronneberger et al., 2015). This U-Net configuration is characterised by a series of levels, where each level has approximately half the resolution of the previous one, and the number of features is doubled. At each level, we apply a sequence of three convolutions, followed by max pooling, and then a transposed convolution for upsampling. After upsampling, three more convolutions are applied at each level. The U-Net used in this study consists of four layers and incorporates the GeLU activation function.